Wendy was four years older than three times the age of Sammy. The sum of their ages is 60. How old is each one.
We know the answers are 14 and 46, we just cant figure out the equation.

Found 2 solutions by scott8148, ptaylor:
Answer by scott8148(6628)

"Wendy was four years older than three times the age of Sammy" ___ w = 3s + 4

"The sum of their ages is 60" ___ w + s = 60

Answer by ptaylor(2198)

OK
Let x=Sammy's age
Three times Sammy's age is 3x
Four years more than three times Sammy's age is 3x+4 and this is Wendy's age, sooo
x+3x+4=60 subtract 4 from each side
4x=56
x=14-----------Sammy's age
3x+4=3*14+4=46---Wendy's age

Hope this helps---ptaylor